# Toward Routing River Water in Land Surface Models with Recurrent Neural Networks

### v1

@@ Abstract
Machine learning is playing an increasing role in hydrology, supplementing or replacing physics-based models. One notable example is the use of recurrent neural networks (RNNs) for forecasting streamflow given observed precipitation and geographic characteristics. Training of such a model over the continental United States has demonstrated that a single set of model parameters can be used across independent catchments, and that RNNs can outperform physics-based models. In this work, we take a next step and study the performance of RNNs for river routing in land surface models (LSMs). Instead of observed precipitation, the LSM-RNN uses instantaneous runoff calculated from physics-based models as an input. We train the model with data from river basins spanning the globe and test it in streamflow hindcasts. The model demonstrates skill at generalization across basins (predicting streamflow in unseen catchments) and across time (predicting streamflow during years not used in training). We compare the predictions from the LSM-RNN to an existing physics-based model calibrated with a similar dataset and find that the LSM-RNN outperforms the physics-based model. Our results give further evidence that RNNs are effective for global streamflow prediction from runoff inputs and motivate the development of complete routing models that can capture nested sub-basis connections.
@@
Within an ESM, a river routing model must demonstrate generalizability across multiple temporal and spatial scales. In time, it should capture the seasonal cycle and the faster surface runoff response to precipitation events. In space, it must display generalizability across basins, one of the main challenges in modern hydrology (Sivapalan et al., 2003; Hrachowitz et al., 2013). River basins, the areas drained by individual rivers and its tributaries, tile the land surface into weakly connected domains, with water transport between basins given by the river streamflow. What we call basin generalizability, known in the hydrology community as regionalization, describes the skill of a model at using the learned characteristics of gauged basins (where streamflow is measured by streamgauges) to predict behavior in other, possibly ungauged, basins, generally in the same region at a sub-continental scale. Several approaches exist to tackle this problem (Prieto et al., 2019). What we call global basin generalizability refers to the same concept, but across all regions, and is also referred to as global-scale regionalization (Beck et al., 2016). In both cases, hydrological models that rely heavily on single-basin calibrations have greater difficulty generalizing due to overfitting to individual basins (Kratzert et al., 2024). Moreover, the vast majority of basins in many regions is ungauged; the generalization of models calibrated with basins from well-represented regions to those with a lack of gauges therefore is especially challenging (Feng et al., 2021).
@@
A physics-based approach to river modeling can be implemented by adapting the shallow-water equations to a 1D setting under various approximations (Li et al., 2013). Two main processes are typically considered: hillslope routing and river channel routing (Mizukami et al., 2016). Hillslope routing describes how water moves across the landscape, considering factors such as topography, soil characteristics, and vegetation. It is the process that addresses the time lag between instantaneously generated runoff on land and the aggregation of runoff at a river outlet, forming the streamflow. This process is typically unresolved in river models used in ESMs and must be parameterized. In contrast, river channel routing describes how water moves from upstream channels to downstream outlets within the river network itself. Both the hillslope and river channel routing processes occur simultaneously within a basin. Physics-based representation of these processes present several advantages. First, the physical equations describing the flow naturally conserve water mass, which is crucial for systems simulated for long periods of time, as is the case in ESMs. Second, the interpretability of the model is straightforward since one knows exactly what physical laws are being used. Finally, physical laws have a long history of success in physical modeling; thus, these approaches are commonly employed in streamflow forecasting models across disciplines. For a more detailed overview of physics-based routing models, see Shaad (2018).
@@
Recently, a class of recurrent deep learning models, referred to as Long-Short-Term-Memory (LSTM) models, have outperformed physics-based models on the rainfall-runoff problem (Kratzert et al., 2019b). In Kratzert et al. (2018), an LSTM was trained to model the entire land hydrology system for the continental United States. The model took observed precipitation as input (along with other dynamic inputs such as near-surface temperature, surface pressure, etc.) and then forecasted streamflow at the outlet of gauged catchments, implicitly modeling snowpack, soil storage, runoff, hillslope routing, and river channel routing. In Kratzert et al. (2019b), attributes such as topography, vegetation, and soil properties from different catchments were added to the model to improve its performance. The model did not explicitly account for routing between basins and treated each catchment independently. Nonetheless, the model showed good performance in regional calibrations, where a single set of parameters is learned using data from multiple catchments at once, employing large-sample hydrology data. Subsequently, predictions were made for the same basins that were used in training. Further studies also showed that these types of models were successful at basin generalization, forecasting basins not included in the training set (Kratzert et al., 2019a). This indicates that information in large-scale hydrological datasets is sufficient for generalization tasks, especially to the common ungauged basins (Nearing et al., 2021). In particular, it also suggests that these models are good candidates for representing processes that are not explicitly resolved in physical models (e.g., because they are below the model’s spatial resolution). Although machine learning (ML) models are not guaranteed to conserve mass a priori (or produce otherwise physical output, such as positive streamflow), constraints can be enforced by adapting the architecture of the network, for example, to be mass conserving (Hoedt et al., 2021).
@@ 0.1 Our contributions
Our goal in this work is to adapt the rainfall-runoff model (an LSTM) of Kratzert et al. (2019b) for use within a global LSM. To do so, we make several changes to the model architecture and training and validation procedures.
@@
We construct a globally consistent dataset to train and validate runoff-driven models. We incorporate runoff variables from reanalysis and use a globally unified system of basin characterizations, to ensure that our routing model can be integrated into an LSM in the future. This also requires addressing consistency between gauged catchments and geographical definitions of the nested sub-basins provided by our base dataset (Lehner et al., 2008). A similar idea was first introduced by the Caravan dataset (Kratzert et al., 2023), using precipitation instead of runoff.
We evaluate the trained LSTM models on LSM-relevant tasks, including generalization across time and basins, using both continental and global training data. We demonstrate good performance of the model and interpret results at a granular level by breaking down skill over different geographical regions.
We compare our generalization experiments with the physics-based LISFLOOD model (Van Der Knijff et al., 2010), which underlies the Global Flood Awareness System (GloFAS), run operationally by the European Copernicus program and at the European Centre for Medium-Range Weather Forecasts (ECMWF). To do so, we use the GloFAS discharge product provided as reanalysis (Harrigan et al., 2020). Results show the RNN approach displays superior performance, and we detail the steps taken to ensure this experiment was as fair as possible.
We note that the terms catchment and basin are often used interchangeably in the hydrology literature; however, in this paper we use “basin” (and sub-basin) for the units of topographical subdivision of the world into drainage areas and “catchment” for the drainage area of specific gauges.
The paper is structured as follows. Section 1 (Methods) presents the various components of our model, including data engineering and the training of the ML model. Section 2 (Results) presents our findings regarding temporal and basin generalization, a comparison to a physics-based model, and an investigation of model performance by basin attributes. Section 3 (Discussion and conclusion) summarizes the main results and outlines future directions.
@@ 1.1 Dataset
The first phase of this project consisted of constructing a consistent dataset that allows world-wide calibrations and simulations. Using global data for training is important as it increases the likelihood of generalization outside the training sample. In an ESM, river routing must be simulated across the entire globe, and not just across basins in the training set. To construct the dataset, we need both forcing data, which vary in time and space, as well as static attributes describing physical characteristics of each basin, which are assumed to only vary in space and which encode how runoff is routed within a basin. We additionally require streamflow data in each of these basins, which is the quantity our model is predicting. As explained in the introduction, we only target within-basin routing (hillslope routing and within-basin channels routing), and not main channel routing between nested sub-basins. This allows us to treat each basin as independent in training and simplifies the training task, at the expense of not making use of the information present in the river network structure. In practical terms, this implies that each catchment represented by a gauge in our data set must have a clear match to a basin.
@@ 1.1.2 Dynamic inputs
A final pre-processing step prior to training and network evaluation is to normalise all input variables, following Kratzert et al. (2022).
@@ 1.1.3 Streamflow
Measurements of streamflow as a function of time were obtained from the Global Runoff Data Centre (GRDC, 2023). To associate the discharge records from the river gauges, identified by latitude and longitude, with their corresponding basins, we employed the ray-casting algorithm to determine which polygon (basin vector) encloses each gauge. We found that in many cases, the gauge catchment area defined by GRDC and the basin area for the corresponding basin in HydroSHEDS were not in agreement. This can occur for two main reasons:
@@ 1.1.4 Dataset summary
The process was applied for all 9 continental areas and for 3 of the 12 levels available in HydroSHEDS (levels 5, 6, and 7). These levels were chosen based on their size (median values of 17,472 km2, 5,318 km, and 1,537 km2, respectively), which is relevant for typical resolutions of climate models. Table 1 summarizes the data. The study covered the time span from 1990 to 2019, with certain gauges exhibiting gaps in discharge data. No gap filling technique was applied, and only original values were retained.
Level | US | Global |
@@
Level 06 | 150 | 443 |
Level 07 | 193 | 660 |
Combined levels | 398 | 1327 |
@@ 1.2 LSTM model
where is a function that is composed of a linear transformations and a dropout layer used to prevent overfitting (Srivastava et al., 2014). A schematic of the network is shown in Figure 3.
LSTMs (Long Short-Term Memory; Hochreiter and Schmidhuber (1997)) are a type of RNN with a specific general structure of the function . A detailed implementation of the LSTM network adapted for rainfall-runoff modeling can be found in Kratzert et al. (2018) and Kratzert et al. (2019b), on which our work is based. The design of the network avoids the vanishing gradient problem that plagues the training procedures for vanilla RNNs, so that optimization of the weights of an LSTM is far more effective. As an RNN, it also allows for the state at previous steps to affect the output at the current step.
Without additional constraints, there is no guarantee that the LSTM model conserves water mass (aside from indirectly, by matching observed streamflow given runoff as input). An important consideration for LSMs/ESMs is how to adapt this model in order to conserve mass, possibly accounting for processes like evaporation from rivers and re-infiltration of water into the soil. Possible approaches include changing the loss function or adapting the neural network design to be mass conserving (Hoedt et al., 2021); we leave such adaptations for future work. Specific details about the hyper-parameters used by our model can be found in Appendix B. More details can also be found in our code (Lima, 2024).
@@ 1.3 Metrics
where and are the predicted and observed streamflow at the the basin outlet for time and is the averaged observed streamflow over all times in the period . In this expression, 1 is a perfect score (), and it gets worse (lower NSE) as the fraction of the mean squared error () normalized by the variance of the streamflow () increases. This normalization implies the NSE to lie in . Note that if a model is predicting only the mean flow at the outlet (i.e., ), we would have an NSE of 0. We will use this value as reference to evaluate good performances (NSE > 0) vs. bad performances (NSE < 0), as suggested in Knoben et al. (2019). The loss function optimized in the training of our models is an adaptation of the NSE. The adaptation and its motivation are explained in Kratzert et al. (2019b).
@@ 2 Results
We present a series of experiments carried out to quantify the performance of the model. We compare the behavior of a model driven with precipitation to that of one driven with runoff, and we compare the behavior of a model trained and tested in the USA with one trained and tested globally. To assess generalizability across basins and times, we experiment with different training/testing/validation splits. The model’s training time varied with the datasets and the longest run took 8h with one V100 GPU. Furthermore, we compare the performance of the model against the physics-based model LISFLOOD (Van Der Knijff et al., 2010), provided by the ECMWF reanalysis data as GloFAS-ERA5 (Harrigan et al., 2020), which uses similar input and calibration data. The hindcasts generated by both models are presented under various conditions, including those where both models produce poor hindcasts, those where the hindcasts are close to the median score, and those where each model demonstrates good performance. Finally, we analyze the performance by continent and other attributes. An analysis by HydroSHEDS levels is provided in Appendix C.
All models shown in this section are based on the basic LSTM architecture. An analysis of mass conservation for our LSTM can be found in Appendix D.

### v3

@@ Abstract
Machine learning is playing an increasing role in hydrology, supplementing or replacing physics-based models. One notable example is the use of recurrent neural networks (RNNs) for forecasting streamflow given observed precipitation and geographic characteristics. Training of such a model over the continental United States (CONUS) demonstrated that a single set of model parameters can be used across independent catchments, and that RNNs can outperform physics-based models. In this work, we take a next step and study the performance of RNNs for river routing in land surface models (LSMs). Instead of observed precipitation, the LSM-RNN uses instantaneous runoff calculated from physics-based models as an input. We train the model with data from river basins spanning the globe and test it in streamflow hindcasts. The model demonstrates skill at generalization across basins (predicting streamflow in catchments not used in training) and across time (predicting streamflow during years not used in training). We compare the predictions from the LSM-RNN to an existing physics-based model calibrated with a similar dataset and find that the LSM-RNN outperforms the physics-based model. Our results show that RNNs are effective for global streamflow prediction from runoff inputs and motivate the development of complete routing models that can capture nested sub-basis connections.
@@
River basins, the areas drained by individual rivers and their tributaries, tile the land surface into weakly connected domains, with water transport between basins given by the river streamflow. Two main processes are typically considered in river modeling: hillslope routing and river channel routing (Mizukami et al., 2016). Hillslope routing describes how water moves across the landscape, considering factors such as topography, soil characteristics, and vegetation. It is the process that addresses the time lag between instantaneously generated runoff on land and the aggregation of runoff at a river channel, contributing to streamflow. This process is typically unresolved in the river models used in LSMs and must be parameterized. In contrast, river channel routing describes how water moves from upstream channels to downstream outlets within the river network itself. Both the hillslope and river channel routing processes occur simultaneously within a basin.
Within an LSM, a river routing model must demonstrate generalizability across multiple temporal and spatial scales. In time, it should capture the seasonal cycle and the faster surface runoff response to precipitation events. In space, it must display generalizability across basins, one of the main challenges in modern hydrology (Sivapalan et al., 2003; Hrachowitz et al., 2013). What we call regional basin generalizability, known in the hydrology community as regionalization, describes the skill of a model at using the learned characteristics of gauged basins (where streamflow is measured by streamgauges) to predict behavior in other, possibly ungauged, basins, generally in the same region at a sub-continental scale. Several approaches exist to tackle this problem (Prieto et al., 2019). What we call global basin generalizability refers to the same concept, but across all regions, and is also referred to as global-scale regionalization (Beck et al., 2016). In both cases, hydrological models that rely heavily on single-basin calibrations have greater difficulty generalizing due to overfitting to individual basins (Kratzert et al., 2024). Moreover, the vast majority of basins in many regions are ungauged; the generalization of models calibrated with basins from well-represented regions to those with a lack of gauges therefore is especially challenging (Feng et al., 2021).
A physics-based approach to river modeling can be implemented considering different processes at different scales. Lohmann et al. (1996) was one of the first river routing schemes for LSMs, using a linear model to account for hillslope routing in each grid and the linearized Saint-Venant equation to model water transport in-between grids. Later models, like the storage-based schemes of Oki and Sud (1998) and Branstetter (2001), used simplified 1D (one dimensional) kinematic wave routing (KWR) equations to route water for both hillslope and river channel routing indistinguishably, but explicitly accounting for spatial distribution of streamflow under various approximations. More advanced approaches (e.g., Ye et al., 2012; Wu et al., 2014), apply 1D KWR explicitly accounting for the differences between hillslope and river channel routing. Moreover, Li et al. (2013) additionally accounted for tributary channels explicitly. More recent models, such as Mizukami et al. (2016), offer flexibility by coupling different possibilities for hillslope and river channel routing in a domain that can be either vector- or grid-based. Physics-based representation of these processes present several advantages. First, the physical equations describing the flow naturally conserve water mass, which is crucial for systems simulated for long periods of time, as is the case in LSMs. Second, the interpretability of the model is straightforward since one knows exactly what physical laws are being used. Finally, physical laws have a long history of success in physical modeling; thus, these approaches are commonly employed in streamflow forecasting models across disciplines. Despite these advantages, however, physical models tend to perform poorly at regional and global basin generalizability, and it has been argued that this is due to challenges in expressing routing processes across scales and locations using simple physical laws (Nearing et al., 2021). For a more detailed overview of physics-based routing models, see Shaad (2018).
@@
Recently, a class of recurrent deep learning models, referred to as Long-Short-Term-Memory (LSTM) models, have outperformed physics-based models on the rainfall-runoff problem (Kratzert et al., 2019b). In Kratzert et al. (2018), an LSTM was trained to model the entire land hydrology system for the CONUS. The model took observed precipitation as input (along with other dynamic inputs such as near-surface temperature, surface pressure, etc.) and then simulated streamflow at the outlet of gauged catchments, implicitly modeling snowpack, soil storage, runoff, hillslope routing, and river channel routing. In Kratzert et al. (2019b), attributes such as topography, vegetation, and soil properties from different catchments were added to the model to improve its performance. The model did not explicitly account for routing between basins and treated each catchment independently. Nonetheless, the model showed good performance in regional calibrations, where a single set of parameters is learned using data from multiple catchments at once. Further studies also showed that these types of models were successful at basin generalization, predicting streamflow on the outlet of catchments not included in the training set (Kratzert et al., 2019a) and more recently, similar models have demonstrated greater global basin generalizability than physics-based models (Nearing et al., 2024). In particular, these results suggest that models can represent the unresolved processes in hillslope and channel routing with accurately. A drawback to using these machine learning (ML) models in LSMs is that they are not guaranteed to conserve mass a priori (or produce otherwise physical output, such as positive streamflow); however, constraints can be enforced by adapting the architecture of the network, for example, to be mass conserving (Hoedt et al., 2021). More importantly, though machine-learning based models have been used for routing between basins in specific regions (Moshe et al., 2020), these types of models have not been used yet for this purpose over entire continents, which would also be a necessary step in order to implement them in LSMs.
@@ 0.1 Our contributions
Our goal in this work is to explore the use of LSTMs for river routing in a global LSM. To do so, we make multiple alterations to the rainfall-runoff model (an LSTM) of Kratzert et al. (2019b), including in model architecture and training and validation procedures, as follows.
@@
We construct a globally consistent dataset to train and validate runoff-driven models. We incorporate runoff variables from reanalysis and use a globally unified system of basin characterizations, to ensure that our routing model can be integrated into an LSM in the future. This also requires addressing consistency between gauged catchments and geographical definitions of the nested sub-basins provided by our base dataset (Lehner et al., 2008). A similar idea was first introduced by the Caravan dataset (Kratzert et al., 2023), using precipitation instead of runoff and catchments instead of globally consistent sub-basins.
We evaluate the trained LSTM models on LSM-relevant tasks, including generalization across time and basins, using both CONUS and global training data. We demonstrate good performance of the model and interpret results at a granular level by breaking down skill over different geographical regions.
We compare our generalization experiments with the physics-based LISFLOOD model (Van Der Knijff et al., 2010), which underlies the Global Flood Awareness System (GloFAS), an operational product and service of the Copernicus Emergency Management Service. To do so, we use the GloFAS discharge product provided as reanalysis. Results show the RNN approach displays superior performance.
We consider the mass conservation properties of the LSTM (in Appendix D).
This work hence represents a first step toward using LSTMs for river routing in LSMs, but we still treat each basin as independent as in Kratzert et al. (2019b). The problem of routing between basins is left for future work.
0.2 Outline and Notation
Regarding notation, we wish to be explicit about the following. We note that the terms catchment and basin are often used interchangeably in the hydrology literature; however, in this paper we use “basin” (and sub-basin) for the units of topographical subdivision of the world into drainage areas and “catchment” for the drainage area of specific gauges. Additionally, we use the term generalization in time as described above (generalizing to time periods unseen in calibration, but only for basins used in calibration) in place of what the hydrology community often refers to as performance in gauged basins. We use the term basin generalization as described above (generalizing to basins not used in calibration) in place of what the hydrology community refers to as regionalization or performance in ungauged basins. This is because whether a basin is gauged or not is a property of the basin, while time and basin generalizability are experimental design concepts. For example, only gauged basins can be used for validating a model, regardless of if that model is calibrated in a “gauged" (generalization in time) or “ungauged" (basin generalization) fashion).
The paper is structured as follows. Section 1 (Methods) presents the various components of our model, including data engineering and the training of the ML model. Section 2 (Results) presents our findings regarding time and basin generalization, a comparison to a physics-based model, and an investigation of model performance by basin attributes. Section 3 (Discussion and conclusion) summarizes the main results and outlines future directions, including applying the LSTM to routing between basins.
@@ 1.1 Dataset
The first phase of this project consisted of constructing a consistent dataset that allows world-wide calibrations and simulations. Using global data for training is important as it increases the likelihood of generalization outside the training sample. In an LSM, river routing must be simulated across the entire globe, and not just across basins in the training set. To construct the dataset, we need both forcing data, which vary in time and space, as well as static attributes describing physical characteristics of each basin, which are assumed to only vary in space and which encode how runoff is routed within a basin. We additionally require streamflow data in each of these basins, which is the quantity our model is predicting. As explained in the introduction, we only target within-basin routing (hillslope routing and within-basin channels routing), and not main channel routing between nested sub-basins. This allows us to treat each basin as independent in training and simplifies the training task, at the expense of not making use of the information present in the river network structure. In practical terms, this implies that each catchment represented by a gauge in our dataset must have a clear match to a basin.
@@ 1.1.2 Dynamic inputs
A final pre-processing step prior to training and network evaluation is to normalize all input variables, following Kratzert et al. (2022).
@@ 1.1.3 Streamflow
Measurements of streamflow as a function of time were obtained from the Global Runoff Data Centre (GRDC, https://portal.grdc.bafg.de/, last accessed 11 September 2024). To associate the discharge records from the river gauges, identified by latitude and longitude, with their corresponding basins, we employed the ray-casting algorithm to determine which polygon (basin vector) encloses each gauge. We found that in many cases, the gauge catchment area defined by GRDC and the basin area for the corresponding basin in HydroSHEDS were not in agreement. This can occur for two main reasons:
@@ 1.1.4 Dataset summary
The process was applied for all 9 continental areas and for 3 of the 12 levels available in HydroSHEDS (levels 5, 6, and 7). These levels were chosen based on their sub-basins area to be relevant for typical resolutions of climate models. Table 1 summarizes the number of catchments in each level both in the CONUS and globally, and reports their areas. The study covered the time span from 1990 to 2019, with certain gauges exhibiting gaps in discharge data. No gap filling technique was applied, and only original values were retained.
| US | Global | |||||
| Level | Number | Median (km2) | Min, Max (km2) | Number | Median (km2) | Min, Max (km2) |
@@
| Level 06 | 150 | 1.1e4 | 1.9e3, 6.2e4 | 443 | 9.9e3 | 1.2e3, 7.0e4 |
| Level 07 | 193 | 4.4e3 | 7.6e2, 2.1e4 | 660 | 3.7e3 | 3.9e2, 5.4e4 |
@@ 1.2 LSTM model
where is a function that is composed of a linear transformation and a dropout layer used to prevent overfitting (Srivastava et al., 2014). A schematic of the network is shown in Figure 3.
LSTMs (Long Short-Term Memory; Hochreiter and Schmidhuber, 1997) are a type of RNN with a specific general structure of the function . A detailed explanation of the LSTM network used in this work can be found in Appendix B. The design of the network avoids the vanishing gradient problem that plagues the training procedures for vanilla RNNs, so that optimization of the weights of an LSTM is far more effective. As an RNN, it also allows for the state at previous steps to affect the output at the current step.
The training procedure seeks to optimize the loss function by varying the free parameters of the LSTM. Our loss function is a modified Nash-Sutcliffe efficiency (NSE), defined as
where and are the predicted and observed streamflow at basin ’s outlet for time in the training set, is the standard deviation of streamflow at basin and is a small number (set to ) for numeric stability. The loss function is averaged by the number of basins in the training set . Moreover, we chose to skip missing values of streamflow to increase the number of valid samples upon which we can compute the loss, which may then vary for each basin. Since there is no factor in the numerator dividing by the number of observations per basin, this indicates that basins with more observations (fewer gaps) are effectively weighted more by the loss function. We refer the reader to Kratzert et al. (2019b) for further discussion of this loss function choice. Note that the NSE is a standard metric for performance used in timeseries prediction; we provide a further description and analysis of the NSE metric in Section 1.3.
To tune the model, we used the Adam optimizer (Kingma and Ba, 2014), with the recommended parameters of , and . Following (Kratzert et al., 2019b), we chose the number of recurrent iterations of the LSTM to be , the number of features in the hidden state to be and the dropout probability of the linear layer in to be . We trained the model for 35 epochs, with a learning rate of for the first 10 epochs, for the 10 following epochs, and for the last 5 epochs. More details can also be found in our code (Lima, 2024).
Without additional constraints, there is no guarantee that the LSTM model conserves water mass (aside from indirectly, by matching observed streamflow given runoff as input). An important consideration for LSMs is how to adapt this model in order to conserve mass, possibly accounting for processes like evaporation from rivers and re-infiltration of water into the soil. Possible approaches include changing the loss function or adapting the neural network design to be mass conserving (Hoedt et al., 2021); we leave such adaptations for future work.
@@ 1.3 Metrics
where and are the predicted and observed streamflow at a basin outlet for time and is the averaged observed streamflow over all valid times in the validation or test set. Note that here we omit the index because the expression is computed for each basin, and not for an ensemble of basins as we did in the expression of the loss function. In this expression, 1 is a perfect score (), and it gets worse (lower NSE) as the fraction of the mean squared error ( normalized by the variance of the streamflow ( increases, where indicates the number of observations for each basin. This normalization implies the NSE to lie in . Note that if a model is predicting only the mean flow at the outlet (i.e., ), we would have an NSE of 0. We will use this value as reference to evaluate good performances (NSE > 0) vs. bad performances (NSE < 0), as suggested in Knoben et al. (2019).
@@ 2 Results
We present a series of experiments carried out to quantify the performance of the model. We compare the behavior of a model driven with precipitation to that of one driven with runoff, and we compare the behavior of a model trained and tested in the USA with one trained and tested globally. To assess generalizability across basins and times, we experiment with different training/testing/validation splits. The model’s training time varied with the datasets and the longest run took a few hours with one V100 GPU. Furthermore, we compare the performance of the model against the physics-based model LISFLOOD (Van Der Knijff et al., 2010), provided by GloFAS v4.0 reanalysis data, which is publicy available in the Copernicus Climate Data Store (https://confluence.ecmwf.int/display/CEMS/GloFAS+v4.0, last accessed 4 September 2024). The hindcasts generated by both models are presented under various conditions, including those where both models produce poor hindcasts, those where the hindcasts are close to the median score, and those where each model demonstrates good performance. An additional point is made regarding possible comparisons with the models presented here and other LSTM models in the literature. Finally, we analyze the performance by continent and other attributes. An analysis by HydroSHEDS levels is provided in Appendix C. All models shown in this section are based on the same LSTM architecture, which is not strictly mass conserving. An analysis of mass conservation for our LSTM can be found in Appendix D.